# Refinements of provability and consistency principles for the second incompleteness theorem

### v1

@@ 1 Introduction
[…]
@@
[…]
Figure 1: Implications between the provability conditions and the versions of G2 for $\Sigma_{1}$ provability predicates
@@ 2 Local derivability conditions
[…]
Here, $\mathrm{Pr}_{\emptyset}(x)$ denotes a standard provability predicate of pure first-order logic in $\mathcal{L}_{A}$. We note that every provability predicate $\mathrm{Pr}_{T}(x)$ satisfies $\mathbf{D1}$. The implications “$\mathbf{D2}\Rightarrow\mathbf{M}$”, “$\mathbf{\Sigma_{1}C}\Rightarrow\mathbf{D3}$”, and “$\mathbf{\Sigma_{1}C}\Rightarrow\mathbf{\Sigma_{1}C}^{-}$” are obvious. Here, “$\mathbf{D2}\Rightarrow\mathbf{M}$” is an abbreviation for the statement “every $\Sigma_{1}$ provability predicate satisfying $\mathbf{D2}$ also satisfies $\mathbf{M}$”. For the sake of simplicity, we adopt this kind of notation throughout the present paper.
@@
    ([17, Proposition 2.4.6 and Remark 2.11]) $\mathbf{M}$ $\mathbf{\Sigma_{1}C}^{-}\Rightarrow\mathbf{\Sigma_{1}C}$.
@@
    ([17, Proposition 2.4.6]) $\mathbf{M}$ $\mathbf{\Sigma_{1}C}\Rightarrow\mathbf{PC}$.
@@
Let $\mathrm{Fml}_{\mathcal{L}_{A}}(x)$ be a natural formula saying that $x$ is a Gödel number of an $\mathcal{L}_{A}$-formula. Let ¬˙$\dot{\neg}\,x$ be a term corresponding to the primitive recursive function calculating the Gödel number of $\neg\,\varphi$ from that of $\varphi$. We may freely use such terms in defining formulas. In [17], the following two consistency statements were distinguished.
@@
    (Löb [22]) $\mathbf{D2}$ $\mathbf{D3}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{L}}_{T}$.
@@
    ([17, Theorems 2.8]) $\mathbf{M}$ $\mathbf{D3}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{H}}_{T}$.
@@
The third clause is a refinement of G2 of Jeroslow [12] and Kreisel and Takeuti [15] stating “$\mathbf{\Sigma_{1}C}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{H}}_{T}$.”
@@ 2.1 Refinements
[…]
@@
The implication “$\mathbf{M}\Rightarrow\mathbf{E}$” and the equivalence “$\mathbf{D2}\Leftrightarrow\mathbf{M}$ $\mathbf{C}$” hold obviously.
@@
    $\mathbf{Ros}$ $\mathbf{C}\Rightarrow T\vdash\mathrm{Con}^{\mathrm{S}}_{T}$.
@@
    $T\vdash\mathrm{Con}^{\mathrm{L}}_{T}$ $\mathbf{E}\Rightarrow\mathbf{Ros}$.
@@
$\mathbf{E}$ $\mathbf{\Sigma_{1}C}^{-}\Rightarrow\mathbf{\Sigma_{1}C}$.
@@
    $\mathbf{E}$ $\mathbf{D3}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{S}}_{T}$.
@@
    $\mathbf{E}$ $\mathbf{C}$ $\mathbf{D3}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{L}}_{T}$.
@@
    $\mathbf{C}$ $\mathbf{\Sigma_{1}C}^{-}\Rightarrow$ non-$\mathbf{Ros}$.
@@
In particular, Corollary 2.9.1 is a refinement of the well-known formulation of G2 presented in Fact 2.4.1, and is of particular interest. Furthermore, we will prove that $\{\mathbf{E},\mathbf{C},\mathbf{D3}\}$ is strictly weaker than $\{\mathbf{D2},\mathbf{D3}\}$ (Theorem 4.2). From Corollary 2.9.2, we have that “$\mathbf{C}$ $\mathbf{\Sigma_{1}C}\Rightarrow\mathrm{Pr}_{T}(x)$ does not satisfy $\mathbf{Ros}$”. Interestingly, this implication is refined by proving the following new version of G2 concerning the newly introduced principle $\mathbf{C}$ and consistency rule $\mathbf{Ros}$.
Theorem 2.10.
$\mathbf{C}$ $\mathbf{D3}\Rightarrow$ non-$\mathbf{Ros}$.
Proof.
Suppose, towards a contradiction, that $\mathrm{Pr}_{T}(x)$ satisfies $\mathbf{C}$, $\mathbf{D3}$, and $\mathbf{Ros}$. By the Fixed Point Theorem, we find a sentence $\varphi$ satisfying
$\mathsf{PA}\vdash\varphi\leftrightarrow\mathrm{Pr}_{T}(\ulcorner\neg\,\mathrm{Pr}_{T}^{+}(\ulcorner\varphi\urcorner)\urcorner),$ (1)
where $\mathrm{Pr}_{T}^{+}(\ulcorner\varphi\urcorner)$ is $\varphi\land\mathrm{Pr}_{T}(\ulcorner\varphi\urcorner)$. From $\mathbf{D3}$, we have
$S\vdash\mathrm{Pr}_{T}(\ulcorner\varphi\urcorner)\to\mathrm{Pr}_{T}(\ulcorner\varphi\urcorner)\land\mathrm{Pr}_{T}(\ulcorner\mathrm{Pr}_{T}(\ulcorner\varphi\urcorner)\urcorner).$
From $\mathbf{C}$, we get
$S\vdash\mathrm{Pr}_{T}(\ulcorner\varphi\urcorner)\to\mathrm{Pr}_{T}(\ulcorner\mathrm{Pr}_{T}^{+}(\ulcorner\varphi\urcorner)\urcorner).$
Furthermore, by applying $\mathbf{C}$ to this with (1),
$S\vdash\mathrm{Pr}_{T}^{+}(\ulcorner\varphi\urcorner)\to\mathrm{Pr}_{T}(\ulcorner\neg\,\mathrm{Pr}_{T}^{+}(\ulcorner\varphi\urcorner)\land\mathrm{Pr}_{T}^{+}(\ulcorner\varphi\urcorner)\urcorner).$
Since $T\vdash\neg\,\mathrm{Pr}_{T}^{+}(\ulcorner\varphi\urcorner)\land\mathrm{Pr}_{T}^{+}(\ulcorner\varphi\urcorner)$, by applying $\mathbf{Ros}$, we obtain
$T\vdash\neg\,\mathrm{Pr}_{T}(\ulcorner\neg\,\mathrm{Pr}_{T}^{+}(\ulcorner\varphi\urcorner)\land\mathrm{Pr}_{T}^{+}(\ulcorner\varphi\urcorner)\urcorner).$
Thus, $T\vdash\neg\,\mathrm{Pr}_{T}^{+}(\ulcorner\varphi\urcorner)$, and hence $S\vdash\mathrm{Pr}_{T}(\ulcorner\neg\,\mathrm{Pr}_{T}^{+}(\ulcorner\varphi\urcorner)\urcorner)$. By (1), $S\vdash\varphi$, and thus $S\vdash\mathrm{Pr}_{T}^{+}(\ulcorner\varphi\urcorner)$. This contradicts the consistency of $T$. ∎
2.2 Löb’s theorem
In this section, we consider Löb’s theorem in two different settings. In usual, Löb’s theorem holds for provability predicates satisfying $\mathbf{D2}$ and $\mathbf{D3}$. Then G2 for $\mathrm{Con}^{\mathrm{L}}_{T}$ is a direct consequence of Löb’s theorem. In Corollary 2.9, we showed that a weaker $\{\mathbf{E},\mathbf{C},\mathbf{D3}\}$ is sufficient for G2 for $\mathrm{Con}^{\mathrm{L}}_{T}$. We show that the set $\{\mathbf{E},\mathbf{C},\mathbf{D3}\}$ is also sufficient for the following variant of Löb’s theorem.
Proposition 2.11.
Suppose that $\mathrm{Pr}_{T}(x)$ satisfies $\mathbf{E}$, $\mathbf{C}$, and $\mathbf{D3}$. Let $\varphi$ be any formula. If $T\vdash\mathrm{Pr}_{T}(\ulcorner\varphi\land\mathrm{Pr}_{T}(\ulcorner\chi\urcorner)\urcorner)\to\varphi$ for all sentences $\chi$, then $T\vdash\varphi$.
Proof.
Suppose that $\mathrm{Pr}_{T}(x)$ satisfies $\mathbf{E}$, $\mathbf{C}$, and $\mathbf{D3}$. Assume that $T\vdash\mathrm{Pr}_{T}(\ulcorner\varphi\land\mathrm{Pr}_{T}(\ulcorner\chi\urcorner)\urcorner)\to\varphi$ holds for all sentences $\chi$. By the Fixed Point Theorem, we find a sentence $\psi$ satisfying the following equivalence:
$\mathsf{PA}\vdash\psi\leftrightarrow(\mathrm{Pr}_{T}(\ulcorner\psi\urcorner)\to\varphi).$
Since $T\vdash(\psi\land\mathrm{Pr}_{T}(\ulcorner\psi\urcorner))\leftrightarrow(\varphi\land\mathrm{Pr}_{T}(\ulcorner\psi\urcorner))$, by applying $\mathbf{E}$, we have $S\vdash\mathrm{Pr}_{T}(\ulcorner\psi\land\mathrm{Pr}_{T}(\ulcorner\psi\urcorner)\urcorner)\leftrightarrow\mathrm{Pr}_{T}(\ulcorner\varphi\land\mathrm{Pr}_{T}(\ulcorner\psi\urcorner)\urcorner)$.
From $\mathbf{D3}$ and $\mathbf{C}$, we get $S\vdash\mathrm{Pr}_{T}(\ulcorner\psi\urcorner)\to\mathrm{Pr}_{T}(\ulcorner\psi\land\mathrm{Pr}_{T}(\ulcorner\psi\urcorner)\urcorner)$, and hence $S\vdash\mathrm{Pr}_{T}(\ulcorner\psi\urcorner)\to\mathrm{Pr}_{T}(\ulcorner\varphi\land\mathrm{Pr}_{T}(\ulcorner\psi\urcorner)\urcorner)$. By the assumption, $T\vdash\mathrm{Pr}_{T}(\ulcorner\varphi\land\mathrm{Pr}_{T}(\ulcorner\psi\urcorner)\urcorner)\to\varphi$, and thus $T\vdash\mathrm{Pr}_{T}(\ulcorner\psi\urcorner)\to\varphi$. By the fixed point equivalence, $T\vdash\psi$ and so $S\vdash\mathrm{Pr}_{T}(\ulcorner\psi\urcorner)$. Therefore, we conclude $T\vdash\varphi$. ∎
Since $\mathsf{PA}\vdash 0=1\leftrightarrow(0=1\land\mathrm{Pr}_{T}(\ulcorner\chi\urcorner))$ for all $\chi$, Corollary 2.9.1 is in fact a consequence of Proposition 2.11.
While the set $\{\mathbf{E},\mathbf{C},\mathbf{D3}\}$ weakens $\{\mathbf{D2},\mathbf{D3}\}$ by replacing $\mathbf{D2}$ with $\{\mathbf{E},\mathbf{C}\}$, we will show that Löb’s theorem still holds when $\mathbf{D3}$ is weakened, as long as $\mathbf{D2}$ is retained. For each provability predicate $\mathrm{Pr}_{T}(x)$, we recursively define $\mathrm{Pr}_{T}^{n}(\ulcorner\varphi\urcorner)$ as follows: $\mathrm{Pr}_{T}^{0}(\ulcorner\varphi\urcorner)$ is $\varphi$; $\mathrm{Pr}_{T}^{n+1}(\ulcorner\varphi\urcorner)$ is $\mathrm{Pr}_{T}(\ulcorner\mathrm{Pr}_{T}^{n}(\ulcorner\varphi\urcorner)\urcorner)$.
@@ 2.2 Löb’s theorem
The usual $\mathbf{D3}$ is $\mathbf{D3}^{1}_{2}$. The condition $\mathbf{D3}^{n}_{m}$ corresponds to the modal axiom $\Box^{n}A\to\Box^{m}A$ which is studied in [13, 20] in the context of non-normal modal logic. The following theorem is a refinement of the usual Löb’s theorem.

### v2

@@ 1 Introduction
The properties of provability predicates have long been analyzed through the normal modal logic $\mathsf{GL}$ (cf. [2, 3, 11, 27]). On the other hand, recent studies [13, 14, 19] have investigated provability predicates via the framework of non-normal modal logic. This approach shows that some principles and rules from non-normal modal logic can be useful in studying the second incompleteness theorem. For example, in the paper [18], the author introduced the schematic consistency statement $\mathrm{Con}^{\mathrm{S}}_{T}=\{\mathrm{Pr}_{T}(\ulcorner\varphi\urcorner)\to\neg\,\mathrm{Pr}_{T}(\ulcorner\neg\,\varphi\urcorner)\mid\varphi$ is a formula}\} that corresponds in modal logic to the axiom D: $\Box A\to\neg\Box\neg A$. It is well known that D is equivalent, in normal modal logic, to the axiom P: $\neg\Box\bot$, which corresponds to the consistency statement $\mathrm{Con}^{\mathrm{L}}_{T}$. However, in non-normal modal logics weaker than the weakest normal modal logic $\mathsf{K}$, the axioms D and P are not equivalent in general. Correspondingly, [18] proved the existence of a Rosser provability predicate for which $\mathrm{Con}^{\mathrm{S}}_{T}$ not is provable, thereby the versions of G2 for $\mathrm{Con}^{\mathrm{L}}_{T}$ and $\mathrm{Con}^{\mathrm{S}}_{T}$ are shown to be distinct.
@@
$\nvdash\mathrm{Con}^{\mathrm{G}}_{T}$$\nvdash\mathrm{Con}^{\mathrm{L}}_{T}$non-$\mathbf{Ros}$$\nvdash\mathrm{Con}^{\mathrm{S}}_{T}$$\nvdash\mathrm{Con}^{\mathrm{H}}_{T}$$\{\mathbf{E},\mathbf{C},\mathbf{D3}\}$$\{\mathbf{E},\mathbf{C},\mathbf{D3}^{n}_{n+k}\}$$\{\mathbf{E},\mathbf{D3}\}$$\{\mathbf{\Sigma_{1}C}\}$$\{\mathbf{C},\mathbf{D3}\}$$\{\mathbf{C},\mathbf{D3}^{n}_{n+k}\}$$\{\mathbf{E}^{\mathrm{U}},\mathbf{CB_{\exists}}\}$$\{\mathbf{D2},\mathbf{D3}\}$$\{\mathbf{M}^{\mathrm{U}}\}$$\{\mathbf{D1^{\mathrm{U}}},\mathbf{D2^{\mathrm{U}}}\}$$\{\mathbf{E}^{\mathrm{U}},\mathbf{CB_{\exists}},\mathbf{C}\}$$\{\mathbf{CB},\mathbf{\Delta_{0}C}^{\mathrm{U}}\}$$\{\mathbf{D2^{\mathrm{G}}},\mathbf{PC}^{\mathrm{G}}\}$
Figure 1: Implications between the derivability conditions and the versions of G2 for $n,k\geq 1$ and $\Sigma_{1}$ provability predicates
@@ 2 Local derivability conditions
There exists a finite subtheory $T_{0}$ of $T$ such that for any $\Sigma_{1}$ sentence $\varphi$, we have $S\vdash\varphi\to\mathrm{Pr}_{T}(\ulcorner\bigwedge T_{0}\to\varphi\urcorner)$. By convention, we regard $\bigwedge\emptyset$ as $0=0$.
Here, $\mathrm{Pr}_{\emptyset}(x)$ denotes a standard provability predicate of pure first-order logic in $\mathcal{L}_{A}$. We note that every provability predicate $\mathrm{Pr}_{T}(x)$ satisfies $\mathbf{D1}$. The implications $\mathbf{D2}\Rightarrow\mathbf{M}$”, $\mathbf{\Sigma_{1}C}\Rightarrow\mathbf{D3}$”, and $\mathbf{\Sigma_{1}C}\Rightarrow\mathbf{\Sigma_{1}C}^{-}$” are obvious. Here, $\mathbf{D2}\Rightarrow\mathbf{M}$” is an abbreviation for the statement “every $\Sigma_{1}$ provability predicate satisfying $\mathbf{D2}$ also satisfies $\mathbf{M}$”. For the sake of simplicity, we adopt this kind of notation throughout the present paper.
@@
    ([17, Proposition 2.4.6 and Remark 2.11]) $\mathbf{M}$ & $\mathbf{\Sigma_{1}C}^{-}\Rightarrow\mathbf{\Sigma_{1}C}$.
@@
    ([17, Proposition 2.4.6]) $\mathbf{M}$ & $\mathbf{\Sigma_{1}C}\Rightarrow\mathbf{PC}$.
@@
Let $\mathrm{Fml}_{\mathcal{L}_{A}}(x)$ be a natural formula saying that $x$ is a Gödel number of an $\mathcal{L}_{A}$-formula. Let $\dot{\neg}\,x$ be a term corresponding to the primitive recursive function calculating the Gödel number of $\neg\,\varphi$ from that of $\varphi$. We may freely use such terms in defining formulas. In [17], the following two consistency statements were distinguished.
@@
    (Löb [22]) $\mathbf{D2}$ & $\mathbf{D3}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{L}}_{T}$.
@@
    ([17, Theorems 2.8]) $\mathbf{M}$ & $\mathbf{D3}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{H}}_{T}$.
@@
The third clause is a refinement of G2 of Jeroslow [12] and Kreisel and Takeuti [15] stating $\mathbf{\Sigma_{1}C}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{H}}_{T}$.”
@@ 2.1 Refinements
As stated in the introduction, in this paper we introduce not only the derivability conditions and consistency statements as above, but also several additional principles and rules corresponding to axioms and rules that have been studied in the context of non-normal modal logic, and we examine their effects. The rule RE: $\dfrac{A\leftrightarrow B}{\Box A\leftrightarrow\Box B}$ and the axiom C: $\Box A\land\Box B\to\Box(A\land B)$ have been extensively studied in non-normal modal logic weaker than $\mathsf{K}$. Moreover, the axiom P: $\neg\,\Box\bot$ corresponds to the consistency statement $\mathrm{Con}^{\mathrm{L}}_{T}$, while the axiom D: $\Box A\to\neg\,\Box\neg\,A$ corresponds to the following schematic consistency statement $\mathrm{Con}^{\mathrm{S}}_{T}$, introduced in [18]. Additionally, we introduce a rule $\mathbf{Ros}$, which corresponds to the modal rule Ros: $\dfrac{\neg\,A}{\neg\,\Box A}$ introduced and investigated in [19].
@@
The implication $\mathbf{M}\Rightarrow\mathbf{E}$” and the equivalence $\mathbf{D2}\Leftrightarrow\mathbf{M}$ & $\mathbf{C}$” hold obviously.
@@
    $\mathbf{Ros}$ & $\mathbf{C}\Rightarrow T\vdash\mathrm{Con}^{\mathrm{S}}_{T}$.
@@
    $T\vdash\mathrm{Con}^{\mathrm{L}}_{T}$ & $\mathbf{E}\Rightarrow\mathbf{Ros}$.
@@
$\mathbf{E}$ & $\mathbf{\Sigma_{1}C}^{-}\Rightarrow\mathbf{\Sigma_{1}C}$.
@@
    $\mathbf{E}$ & $\mathbf{D3}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{S}}_{T}$.
@@
    $\mathbf{E}$ & $\mathbf{C}$ & $\mathbf{D3}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{L}}_{T}$.
@@
    $\mathbf{C}$ & $\mathbf{\Sigma_{1}C}^{-}\Rightarrow$ non-$\mathbf{Ros}$.
@@
In particular, Corollary 2.9.1 is a refinement of the well-known formulation of G2 presented in Fact 2.4.1, and is of particular interest. Furthermore, we will prove that $\{\mathbf{E},\mathbf{C},\mathbf{D3}\}$ is strictly weaker than $\{\mathbf{D2},\mathbf{D3}\}$ (Theorem 4.2).
From Corollary 2.9.2, we have that $\mathbf{C}$ & $\mathbf{\Sigma_{1}C}\Rightarrow\mathrm{Pr}_{T}(x)$ does not satisfy $\mathbf{Ros}$”. Interestingly, this implication is refined by proving the following new version of G2 concerning the newly introduced principles $\mathbf{C}$, $\mathbf{D3}^{n}_{m}$, and consistency rule $\mathbf{Ros}$. For each provability predicate $\mathrm{Pr}_{T}(x)$, we recursively define $\mathrm{Pr}_{T}^{n}(\ulcorner\varphi\urcorner)$ as follows: $\mathrm{Pr}_{T}^{0}(\ulcorner\varphi\urcorner)$ is $\varphi$; $\mathrm{Pr}_{T}^{n+1}(\ulcorner\varphi\urcorner)$ is $\mathrm{Pr}_{T}(\ulcorner\mathrm{Pr}_{T}^{n}(\ulcorner\varphi\urcorner)\urcorner)$.
@@
The usual $\mathbf{D3}$ is $\mathbf{D3}^{1}_{2}$. The condition $\mathbf{D3}^{n}_{m}$ corresponds to the modal axiom $\Box^{n}A\to\Box^{m}A$ which is studied in [13, 20] in the context of non-normal modal logic.
@@
Let $n,k\geq 1$. Then, $\mathbf{C}$ & $\mathbf{D3}^{n}_{n+k}\Rightarrow$ non-$\mathbf{Ros}$.
Proof.
Suppose, towards a contradiction, that $\mathrm{Pr}_{T}(x)$ satisfies $\mathbf{C}$, $\mathbf{D3}^{n}_{n+k}$, and $\mathbf{Ros}$. By the Fixed Point Theorem, we find a sentence $\varphi$ satisfying
$\mathsf{PA}\vdash\varphi\leftrightarrow\mathrm{Pr}_{T}(\ulcorner\neg\,\xi\urcorner),$ (1)
where $\xi$ is the sentence
$\bigwedge_{0\leq i\leq n+k-1}\mathrm{Pr}_{T}^{i}(\ulcorner\varphi\urcorner).$
From $\mathbf{D3}^{n}_{n+k}$ and (1), we get
$S\vdash\xi$ $\to\varphi\land\bigwedge_{1\leq i\leq n+k-1}\mathrm{Pr}_{T}^{i}(\ulcorner\varphi\urcorner)$
$\to\mathrm{Pr}_{T}(\ulcorner\neg\,\xi\urcorner)\land\bigwedge_{1\leq i\leq n+k}\mathrm{Pr}_{T}^{i}(\ulcorner\varphi\urcorner).$
By the $n+k$ times applications of $\mathbf{C}$, we have
$S\vdash\xi\to\mathrm{Pr}_{T}\left(\ulcorner\neg\,\xi\land\xi\urcorner\right).$
Since $T\vdash\neg\,(\neg\,\xi\land\xi)$, by applying $\mathbf{Ros}$, we obtain $T\vdash\neg\,\mathrm{Pr}_{T}(\ulcorner\neg\,\xi\land\xi\urcorner)$. Thus, $T\vdash\neg\,\xi$, and hence $S\vdash\mathrm{Pr}_{T}(\ulcorner\neg\,\xi\urcorner)$. By (1), $S\vdash\varphi$. By applying $\mathbf{D1}$, we have $S\vdash\bigwedge_{0\leq i\leq n+k-1}\mathrm{Pr}_{T}^{i}(\ulcorner\varphi\urcorner)$. So, $T\vdash\xi$. This contradicts the consistency of $T$. ∎
The following corollary is a refinement of Corollary 2.9, which follows from the combination of Proposition 2.6.5 and Theorem 2.11.
Corollary 2.12.
Let $n,k\geq 1$. Then, $\mathbf{E}$ & $\mathbf{C}$ & $\mathbf{D3}^{n}_{n+k}\Rightarrow T\nvdash\mathrm{Con}^{\mathrm{L}}_{T}$.
2.2 Löb’s theorem
In usual, Löb’s theorem holds for provability predicates satisfying $\mathbf{D2}$ and $\mathbf{D3}$. Then G2 for $\mathrm{Con}^{\mathrm{L}}_{T}$ is a direct consequence of Löb’s theorem. In Corollary 2.12, we showed that a weaker $\{\mathbf{E},\mathbf{C},\mathbf{D3}^{n}_{n+k}\}$ with $n,k\geq 1$ is sufficient for G2 for $\mathrm{Con}^{\mathrm{L}}_{T}$. We show that the set $\{\mathbf{E},\mathbf{C},\mathbf{D3}^{n}_{n+k}\}$ is also sufficient for the following weak variant of Löb’s theorem.
Theorem 2.13.
Let $n,k\geq 1$. Suppose that $\mathrm{Pr}_{T}(x)$ satisfies $\mathbf{E}$, $\mathbf{C}$, and $\mathbf{D3}^{n}_{n+k}$. Let $\varphi$ be any formula. If $T\vdash\mathrm{Pr}_{T}(\ulcorner\varphi\land\bigwedge_{1\leq i\leq n+k-1}\mathrm{Pr}_{T}^{i}(\ulcorner\chi\urcorner)\urcorner)\to\varphi$ holds for all sentences $\chi$, then $T\vdash\varphi$.
Proof.
Suppose that $\mathrm{Pr}_{T}(x)$ satisfies $\mathbf{E}$, $\mathbf{C}$, and $\mathbf{D3}^{n}_{n+k}$. Assume that $T\vdash\mathrm{Pr}_{T}(\ulcorner\varphi\land\bigwedge_{1\leq i\leq n+k-1}\mathrm{Pr}_{T}^{i}(\ulcorner\chi\urcorner)\urcorner)\to\varphi$ holds for all sentences $\chi$. By the Fixed Point Theorem, we find a sentence $\psi$ satisfying the following equivalence:
$\mathsf{PA}\vdash\psi\leftrightarrow(\xi\to\varphi),$
where $\xi$ is the sentence
$\bigwedge_{1\leq i\leq n+k-1}\mathrm{Pr}_{T}^{i}(\ulcorner\psi\urcorner).$
Since $T\vdash(\psi\land\xi)\leftrightarrow(\varphi\land\xi)$, by applying $\mathbf{E}$, we have
$S\vdash\mathrm{Pr}_{T}(\ulcorner\psi\land\xi\urcorner)\leftrightarrow\mathrm{Pr}_{T}(\ulcorner\varphi\land\xi\urcorner).$ (2)
From $\mathbf{D3}^{n}_{n+k}$,
$S\vdash\xi\to\bigwedge_{1\leq i\leq n+k}\mathrm{Pr}_{T}^{i}(\ulcorner\psi\urcorner).$
By the n+k1n+k-1 times applications of $\mathbf{C}$, we get
$S\vdash\xi\to\mathrm{Pr}_{T}\left(\ulcorner\bigwedge_{0\leq i\leq n+k-1}\mathrm{Pr}_{T}^{i}(\ulcorner\psi\urcorner)\urcorner\right),$
and equivalently, $S\vdash\xi\to\mathrm{Pr}_{T}(\ulcorner\psi\land\xi\urcorner)$. It follows from (2) that $S\vdash\xi\to\mathrm{Pr}_{T}(\ulcorner\varphi\land\xi\urcorner)$. By the assumption, $T\vdash\mathrm{Pr}_{T}(\ulcorner\varphi\land\xi\urcorner)\to\varphi$, and thus $T\vdash\xi\to\varphi$. By the fixed point equivalence, $T\vdash\psi$. So, $S\vdash\bigwedge_{1\leq i\leq n+k-1}\mathrm{Pr}_{T}^{i}(\ulcorner\psi\urcorner)$, that is, $S\vdash\xi$. Therefore, we conclude $T\vdash\varphi$. ∎
Since $\mathsf{PA}\vdash 0=1\leftrightarrow(0=1\land\bigwedge_{1\leq i\leq n+k-1}\mathrm{Pr}_{T}^{i}(\ulcorner\chi\urcorner))$ for all $\chi$, Corollary 2.12 is in fact a consequence of Theorem 2.13. The following corollary is a refinement of the usual Löb’s theorem.
Corollary 2.14.